OpenStudy (anonymous):

Set up integrals for both orders of integration, and use the more convenient order to evaluate the integral over the region R.

OpenStudy (anonymous):

\[\int\limits_{R}^{}\int\limits_{}^{}xe^ydA\]

OpenStudy (anonymous):

R:Triangle bounded by y= 4-x , y = 0 , x= 0

OpenStudy (anonymous):

I'm guessing in order to find the other limit I would have to graph xe^y and find out where they both intersect right?